OpenStudy (anonymous):

An object is at rest on the plane. Three forces, V, W, X are acting on the object. If V=9 I − 5 J, W=1 I − 3 J, then X must be ??? I+ ??? J

OpenStudy (anonymous):

v+w+x=0 Start with this

OpenStudy (anonymous):

-(V+W)=-(10 I-8J)=-10I+8J

OpenStudy (anonymous):

hey thanks :D very helpful !